OpenStudy (anonymous):

Explain how to solve: a nurse earning an annual salary of $39,235 gets a 4% raise. What is the amount of the raise? I got $1569.4.....im confused shouldn't it be bigger than her starting salary? and then you would subtract the two?

OpenStudy (anonymous):

you should multiply the salary by 1.04 to get the new salary. Multiplying by just 0.04 gives you the hike.

OpenStudy (anonymous):

where did 1.04 come from?

OpenStudy (anonymous):

4 % hike added to old salaray which is 100% of itself, 100% can be written as 1 when you write it in fraction form

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Amt. of the raise = 4/100 * 39235 = $1569.4 The answer you got is right.

OpenStudy (anonymous):

amount of raise=4/100*39235

OpenStudy (anonymous):

i should not be bigger than her starting salary because they are asking amount the amount raised,not the total amount we get after raising

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

oh yes ! :P